# Single-Round Scalable Analytic Federated Learning

### v1

@@ 1 Introduction
Federated Learning (FL) enables multiple clients or devices to collaboratively train a shared model without exposing their private data. Instead of centralizing data, clients perform local updates and periodically communicate model parameters to a server, which aggregates them into a global model [McMahan2017FedAvg]. While conceptually appealing, conventional FL frameworks require many communication rounds—often hundreds or thousands—for a model to converge. In practical deployments, clients can operate at different speeds, disconnect intermittently, or fail mid-training, creating stragglers and asynchronous updates. Such instability causes training to progress unevenly, and the global model may take days or weeks to reach convergence, severely limiting FL’s real-world scalability.
@@
Beyond communication inefficiency, a deeper issue lies in statistical heterogeneity across clients. In real FL systems, local data distributions often differ sharply—for instance, users capture different visual styles, hospitals record different patient populations, or sensors observe non-overlapping environments. This non-IID nature of the data means that each client’s gradient direction diverges from the global optimum, degrading performance and convergence stability. Existing methods attempt to address this through various regularizers, dynamic aggregation schemes, and using pre-trained models for initialization and distilation [Li2020FedProx, Wang2020FedNova, Li2021MOON, Acar2021FedDyn], but these stuggles with non-IID settings.
@@
To overcome these limitations, recent work proposed Analytic Federated Learning (AFL) [He2025AFL], which formulates the FL problem in closed form. AFL leverages a pre-trained backbone to extract embeddings on each client, and trains a linear regression head analytically in only one communication round. Its analytic aggregation law guarantees invariance to both data partitioning and client count, enabling the global solution to remain identical to centralized training regardless of heterogeneity. As a result, AFL achieves higher accuracy than conventional iterative FL methods under highly non-IID conditions, while requiring only a single communication round instead of hundreds. Despite these appealing properties, AFL remains constrained by its linear model structure, which limits representational capacity and the ability to capture nonlinear feature interactions.
@@
More recently, DeepAFL [deepafl2025] proposed a layer-wise analytic training scheme that extends AFL into deeper architectures. DeepAFL retains AFL’s invariance property, but trades communication efficiency for greater accuracy. Each analytic layer requires a separate aggregation round, increasing synchronization overhead and deviating from AFL’s single-pass analytic design. Consequently, DeepAFL achieves higher accuracy than AFL on non-IID data but at the cost of multiple communication rounds.
@@
In this work, we propose SAFLeSparse Analytic Federated Learning with nonlinear embeddings — a framework that retains AFL’s single-round analytic formulation while significantly enhancing model expressivity. SAFLe introduces a deterministic nonlinear transformation pipeline composed of three stages: feature bucketing, shuffling and grouping and sparse embeddings. We prove that this nonlinear transformation pipeline can be reformulated as an equivalent analytic regression problem, preserving AFL’s closed-form training and invariance properties.
@@ 2.1 Federated Learning
Conventional iterative FL frameworks, pioneered by FedAvg [McMahan2017FedAvg], train a global model through multiple rounds of communication. In this paradigm, clients train local models, and a central server periodically averages their weights. While effective, this iterative approach faces significant challenges with non-independently and identically distributed (non-IID) data. A large body of research has been proposed to address this. Methods like FedProx [Li2020FedProx] add a proximal term to the local objective to restrict updates, while FedNova [Wang2020FedNova] normalizes local updates before averaging. Other major approaches include Model-Contrastive Federated Learning (MOON) [Li2021MOON], which utilizes contrastive learning; FedDyn [Acar2021FedDyn], which uses dynamic regularization; and various knowledge distillation techniques like FedNTD [Lee2022FedNTD], FedGen [Zhu2021FedGen], and FedDisco [Ye2023FedDisco]. Despite their different strategies, all these methods are fundamentally gradient-based and iterative, incurring substantial communication rounds.
To address the high communication overhead, One-Shot Federated Learning (OFL) methods have been proposed [Allouah2024Revisiting]. These methods aim to train a final model with only a single communication round, where clients train models locally and send them to the server once. Common OFL strategies involve ensembling the local models on the server [Allouah2024Revisiting], often using knowledge distillation (KD) to create a single, unified model [Dai2024CoBoosting, Wu2022FedKD]. Methods like FedCVAE-Ens use generative models to facilitate this one-shot distillation [Heinbaugh2023FedCVAE]. However, this efficiency comes at a well-documented cost: OFL methods typically suffer a significant drop in accuracy compared to their iterative counterparts, especially in settings with high data heterogeneity [Allouah2024Revisiting].
@@ 2.2 Analytic Federated Learning (AFL)
Analytic Federated Learning (AFL) [He2025AFL] was proposed to solve both problems. It is a gradient-free, single-round framework that uses a pre-trained frozen backbone. On each client $k$, a linear classification head is trained by solving a simple least-squares (LS) problem:
@@ 2.3 Deep Analytic Federated Learning (DeepAFL)
DeepAFL [deepafl2025] was introduced to address AFL’s linearity constraint by building a deeper, non-linear analytic model inspired by ResNet, that achieves non-linearity by using layers with random weights.
@@ 3.1 Intuition
This architecture is conceptually similar to a Mixture of Experts (MoE) model [jacobs1991adaptive], as it employs a set of ”experts” (our $E$ embedding layers) to learn specialized functions over different parts of the input space. However, SAFLe differs in two crucial ways: its routing mechanism and its activation pattern. First, instead of a learned gating network that dynamically routes an input, SAFLe uses our deterministic bucketing-and-shuffling pipeline as a fixed pre-router. Second, rather than the sparse activation of a standard MoE, SAFLe uses dense activation: all $E$ experts are activated for every input.
@@ 3.3 Analytic Formulation and Aggregation
(Federated Analytic Solution). By extending the RI-AA Law from [He2025AFL], the global model $W_{global}$ can be solved analytically in a single communication round.
@@
Following the RI process [He2025AFL], each client $k$ computes and transmits two matrices based on its local high-dimensional features $\Phi_{k}$:
@@ 4.1 Experimental Setup
Datasets & Settings. We conduct our experiments on three prominent federated learning benchmark datasets: CIFAR-10 [Krizhevsky2009Learning], CIFAR-100 [Krizhevsky2009Learning], and Tiny-ImageNet [Le2015Tiny]. To simulate statistical heterogeneity, we follow the standard setup used in AFL [He2025AFL] and DeepAFL [deepafl2025] and employ two common non-IID partitioning strategies [Lin2020Ensemble]: Non-IID-1 (LDA), where data is partitioned using a Latent Dirichlet Allocation (LDA) process controlled by the parameter $\alpha$, and Non-IID-2 (Sharding), where data is sorted by label and divided into shards, with each client receiving $s$ shards. In both settings, smaller values of $\alpha$ or $s$ indicate a higher degree of data heterogeneity. For CIFAR-1.0, we test with $\alpha\in\{0.1,0.05\}$ and $s\in\{2,4\}$. For the more complex CIFAR-100 and Tiny-ImageNet datasets, we use $\alpha\in\{0.1,0.01\}$ and $s\in\{5,10\}$. Unless otherwise specified, all experiments use a total of $K=100$ clients.
Baselines & Metrics. We compare our proposed SAFLe against two categories of federated learning methods. The first category, Iterative Baselines, includes five widely-used gradient-based methods: FedAvg [McMahan2017FedAvg], FedProx [Li2020FedProx], MOON [Li2021MOON], FedDyn [Acar2021FedDyn], and FedNTD [Lee2022FedNTD]. The second, Analytic Baselines, includes the two most relevant analytic methods: AFL [He2025AFL], the single-round linear analytic model, and DeepAFL [deepafl2025], the state-of-the-art multi-round analytic model. For a fair and direct comparison, all methods (both iterative and analytic) are initialized with the same ResNet-18 backbone pre-trained on ImageNet. The results for all baseline methods are directly obtained from the established benchmark provided in [He2025AFL, deepafl2025]. We evaluate all approaches on Top-1 Accuracy (%) to measure model performance, Communication Rounds to measure synchronization overhead, and Communication Cost (MB) to measure total data transmission per client.
@@ 4.2 Experimental Results
Accuracy Comparisons The main accuracy results, presented in Table 1, show that SAFLe consistently and significantly outperforms all iterative and analytic baselines across all datasets and non-IID settings. On CIFAR-10, SAFLe achieves 90.73% accuracy, surpassing the next-best analytic method, DeepAFL [deepafl2025] (86.43%), by over 4.3% and the original AFL [He2025AFL] (80.75%) by a 10% margin. On CIFAR-100, SAFLe (70.61%) again leads DeepAFL (66.98%) and AFL (58.56%), and on Tiny-ImageNet, our method (64.58%) continues to outperform DeepAFL (62.35%) and AFL (54.67%). This demonstrates the superior representational power of our proposed non-linear analytic architecture, which achieves higher accuracy than even the multi-round DeepAFL while maintaining single-round efficiency.
@@
Invariance Analysis Table 1 empirically shows the core theoretical benefit of our analytic formulation: invariance to data heterogeneity. While all iterative gradient-based methods suffer severe performance degradation as the data becomes more non-IID (e.g., $\alpha=0.01$ or $s=5$), our method does not. For instance, on CIFAR-100, FedAvg’s accuracy plummets from 56.62% ($\alpha=0.1$) to 32.99% ($\alpha=0.01$), a drop of over 23.6%. In stark contrast, all analytic methods are completely unaffected by the data distribution. Our proposed SAFLe achieves an identical 70.61% accuracy across all four non-IID settings on CIFAR-100, just as AFL (58.56%) and DeepAFL (66.98%) maintain their respective scores. This invariance also extends to the number of clients, as shown in Figure 5. While FedAvg’s performance clearly declines as $K$ increases from 100 to 1000, SAFLe’s accuracy remains constant, matching the invariance of AFL. This experimentally showcases SAFLe successful extension of the mathematical invariance of AFL to a more expressive non-linear model.
Communication Rounds Efficiency DeepAFL [deepafl2025] scales AFL by adding depth (i.e., more layers), but at the direct trade-off of more communication rounds. As seen in Figures 4 and 2, DeepAFL must introduce two new communication rounds for every additional layer, requiring 41 rounds for a $T=20$ model. Our method, SAFLe, scales by adding width (i.e., more experts $E$), allowing us to scale the model’s expressivity while remaining a single-shot solution. SAFLe’s expressivity is decoupled from its number of communication rounds, which is always one round.
@@
Model Size Comparison. We note that the additional parameters introduced by our SAFLe architecture—the sparse embedding layers—are negligible in size compared to the backbone. In all configurations, the total size of these embeddings is less than 0.3MB, which is insignificant compared to the $\approx 44$MB ResNet-18. This ensures a fair comparison, as it demonstrates that our method’s significant accuracy improvements are not attributable to an unfair advantage in total model parameters. While SAFLe’s non-linear design *does* scale model capacity, which is the source of its improved expressivity, it does so with a fractional parameter increase so small that it does not to provide any meaningful unfair advantage to the standard iterative, gradient-based baselines. As a reference, this additional size is equivalent to adding approximately 146 neurons to the final hidden layer head of the other SGD baseline methods, which produces no meaningful change in accuracy.
@@
Training Time Comparison. All experiments were run on an NVIDIA RTX 4090 GPU, matching the hardware used in the prior analytic works [He2025AFL, deepafl2025]. DeepAFL (Figure 3 in [deepafl2025]) reports total training times of 130.7s, 145.7s, and 190.0s for its $T=5,10,20$ models on Tiny-ImageNet. To reach the same accuracy levels (60.3%, 61.4%, 62.4%), our SAFLe models are computationally faster, taking 114s, 128s, and 143s, respectively.
@@
Ablation: Bucketing Strategy The choice of bucketing function, which discretizes the continuous backbone features, is critical. We compare three strategies in Table 2: Standard (Integer), One-Hot, and Binary Overlapping. The results show that Binary Overlapping is the best-performing method, while standard Integer quantization is by far the worst. We attribute this to how each method handles local smoothness. Integer quantization creates a ”cliff effect”: a small change in an input feature that crosses a bin boundary results in a completely different index, leading to a drastic change in the looked-up embeddings. This prevents the model from generalizing to similar, nearby inputs. One-Hot binning is slightly better but still maps similar inputs to disjoint representations. The Binary Overlapping method excels because, by design, close input values share similar bit patterns. This allows the model to index and learn similar patterns on the embeddings for similar inputs, leading to significantly better generalization.
@@
Ablation: Different Backbones. To demonstrate that our method’s advantages are general and not specific to one backbone, we follow the ablation protocol from the AFL paper [He2025AFL] and compare SAFLe against AFL using the same three pre-trained and frozen backbones: ResNet-18 [He2016ResNet], VGG11 [Simonyan2015VGG], and ViT-B-16 [Dosovitskiy2021ViT]. The DeepAFL paper [deepafl2025] does not include results for this specific experiment, and their code is not yet publicly available. The results in Table 3 show that SAFLe consistently outperforms AFL across all backbones and datasets. While a stronger backbone like ViT-B-16 significantly increases the performance of both methods [He2025AFL], our non-linear analytic head still provides a clear and consistent accuracy gain over AFL’s linear head. For instance, on CIFAR-100, SAFLe improves AFL’s accuracy from 58.56% to 70.61% with ResNet-18, and from 75.45% to 77.83% with ViT-B-16. This confirms that the expressive power of our non-linear architecture provides a robust benefit, independent of the initial feature quality.

### v2

@@ 1 Introduction
Federated Learning (FL) enables multiple clients or devices to collaboratively train a shared model without exposing their private data. Instead of centralizing data, clients perform local updates and periodically communicate model parameters to a server, which aggregates them into a global model [21]. While conceptually appealing, conventional FL frameworks require many communication rounds—often hundreds or thousands—for a model to converge. In practical deployments, clients can operate at different speeds, disconnect intermittently, or fail mid-training, creating stragglers and asynchronous updates. Such instability causes training to progress unevenly, and the global model may take days or weeks to reach convergence, severely limiting FL’s real-world scalability.
@@
Beyond communication inefficiency, a deeper issue lies in statistical heterogeneity across clients. In real FL systems, local data distributions often differ sharply—for instance, users capture different visual styles, hospitals record different patient populations, or sensors observe non-overlapping environments. This non-IID nature of the data means that each client’s gradient direction diverges from the global optimum, degrading performance and convergence stability. Existing methods attempt to address this through various regularizers, dynamic aggregation schemes, personalization strategies, architectural adaptations, and the use of pre-trained models for initialization and distillation [19, 27, 18, 1, 31, 23, 17, 5], but these methods still struggle under strong non-IID settings.
@@
To overcome these limitations, recent work proposed Analytic Federated Learning (AFL) [11], which formulates the FL problem in closed form. AFL leverages a pre-trained backbone to extract embeddings on each client, and trains a linear regression head analytically in only one communication round. Its analytic aggregation law guarantees invariance to both data partitioning and client count, enabling the global solution to remain identical to centralized training regardless of heterogeneity. As a result, AFL achieves higher accuracy than conventional iterative FL methods under highly non-IID conditions, while requiring only a single communication round instead of hundreds. Despite these appealing properties, AFL remains constrained by its linear model structure, which limits representational capacity and the ability to capture nonlinear feature interactions.
@@
More recently, DeepAFL [25] proposed a layer-wise analytic training scheme that extends AFL into deeper architectures. DeepAFL retains AFL’s invariance property, but trades communication efficiency for greater accuracy. Each analytic layer requires a separate aggregation round, increasing synchronization overhead and deviating from AFL’s single-pass analytic design. Consequently, DeepAFL achieves higher accuracy than AFL on non-IID data but at the cost of multiple communication rounds.
In parallel, recent one-shot federated learning methods have explored ensembling, synthetic data distillation, and pre-trained feature statistics to improve performance under severe heterogeneity [4, 32, 9]. However, these methods still generally lack the exact centralized-equivalence and heterogeneity-invariance guarantees provided by analytic formulations.
@@
In this work, we propose SAFLeSparse Analytic Federated Learning with nonlinear embeddings — a framework that retains AFL’s single-round analytic formulation while significantly enhancing model expressivity. SAFLe introduces a deterministic nonlinear transformation pipeline composed of three stages: feature bucketing, shuffling and grouping, and sparse embeddings. We prove that this nonlinear transformation pipeline can be reformulated as an equivalent analytic regression problem, preserving AFL’s closed-form training and invariance properties.
@@ 2.1 Federated Learning
Conventional iterative FL methods, pioneered by FedAvg [21], train a global model through many communication rounds by alternating local client updates with server-side aggregation. While effective, this paradigm is highly vulnerable to statistical heterogeneity across clients. Numerous methods have been proposed to mitigate this issue, including regularization-based approaches such as FedProx [19] and FedDyn [1], normalized aggregation in FedNova [27], contrastive learning in MOON [18], feature distillation in FedFed [31], and personalization with pre-trained models in FedL2P [17]. Recent work has also highlighted the importance of architectural design in handling heterogeneity for visual FL [23]. Despite these advances, these methods remain fundamentally gradient-based and iterative, incurring substantial communication overhead.
To reduce communication, One-Shot Federated Learning (OFL) aims to complete training in a single round [4]. Existing OFL methods typically rely on server-side ensembling [4], knowledge distillation [7, 28], or more recent synthetic distillation and global feature statistics, as in FedSD2C [32] and FedCGS [9]. However, OFL methods generally still suffer a noticeable accuracy gap relative to iterative FL, especially under strong heterogeneity [4].
@@ 2.2 Analytic Federated Learning (AFL)
Analytic Federated Learning (AFL) [11] was proposed to solve both problems. It is a gradient-free, single-round framework that uses a pre-trained frozen backbone. On each client $k$, a linear classification head is trained by solving a simple least-squares (LS) problem:
@@ 2.3 Deep Analytic Federated Learning (DeepAFL)
DeepAFL [25] was introduced to address AFL’s linearity constraint by building a deeper, non-linear analytic model inspired by ResNet, that achieves non-linearity by using layers with random weights.
@@
2.4 Federated WiSARD Models
A related direction [22] explores federated aggregation with Weightless Neural Networks (WNNs) [3, 2], using counter-based RAM memories to combine local knowledge across clients in a way that is invariant to how the data is partitioned. However, unlike analytic federated learning methods, these approaches do not derive the global model from a closed-form analytical solution; instead, they rely on heuristic-based learning and approximate memory mechanisms, which leads to lower accuracy. As a result, although they retain heterogeneity-invariant aggregation through counter-based RAM structures, they lack the precision and performance of analytically derived federated solutions.
@@ 3.1 Intuition
This architecture is conceptually similar to a Mixture of Experts (MoE) model [13], as it employs a set of ”experts” (our $E$ embedding layers) to learn specialized functions over different parts of the input space. However, SAFLe differs in two crucial ways: its routing mechanism and its activation pattern. First, instead of a learned gating network that dynamically routes an input, SAFLe uses our deterministic bucketing-and-shuffling pipeline as a fixed pre-router. Second, rather than the sparse activation of a standard MoE, SAFLe uses dense activation: all $E$ experts are activated for every input.
@@ 3.3 Analytic Formulation and Aggregation
(Federated Analytic Solution). By extending the RI-AA Law from [11], the global model $W_{global}$ can be solved analytically in a single communication round.
@@
Following the RI process [11], each client $k$ computes and transmits two matrices based on its local high-dimensional features $\Phi_{k}$:
@@ 4.1 Experimental Setup
Datasets & Settings. We conduct our experiments on three prominent federated learning benchmark datasets: CIFAR-10 [14], CIFAR-100 [14], and Tiny-ImageNet [15]. To simulate statistical heterogeneity, we follow the standard setup used in AFL [11] and DeepAFL [25] and employ two common non-IID partitioning strategies [20]: Non-IID-1 (LDA), where data is partitioned using a Latent Dirichlet Allocation (LDA) process controlled by the parameter $\alpha$, and Non-IID-2 (Sharding), where data is sorted by label and divided into shards, with each client receiving $s$ shards. In both settings, smaller values of $\alpha$ or $s$ indicate a higher degree of data heterogeneity. For CIFAR-10, we test with $\alpha\in\{0.1,0.05\}$ and $s\in\{2,4\}$. For the more complex CIFAR-100 and Tiny-ImageNet datasets, we use $\alpha\in\{0.1,0.01\}$ and $s\in\{5,10\}$. Unless otherwise specified, all experiments use a total of $K=100$ clients.
Baselines & Metrics. We compare our proposed SAFLe against two categories of federated learning methods. The first category, Iterative Baselines, includes five widely-used gradient-based methods: FedAvg [21], FedProx [19], MOON [18], FedDyn [1], and FedNTD [16]. The second, Analytic Baselines, includes the two most relevant analytic methods: AFL [11], the single-round linear analytic model, and DeepAFL [25], the state-of-the-art multi-round analytic model. For a fair and direct comparison, all methods (both iterative and analytic) are initialized with the same ResNet-18 backbone pre-trained on ImageNet. The results for all baseline methods are directly obtained from the established benchmark provided in [11, 25]. We evaluate all approaches on Top-1 Accuracy (%) to measure model performance, Communication Rounds to measure synchronization overhead, and Communication Cost (MB) to measure total data transmission per client.
@@ 4.2 Experimental Results
Accuracy Comparisons The main accuracy results, presented in Table 1, show that SAFLe consistently and significantly outperforms all iterative and analytic baselines across all datasets and non-IID settings. On CIFAR-10, SAFLe achieves 90.73% accuracy, surpassing the next-best analytic method, DeepAFL [25] (86.43%), by over 4.3% and the original AFL [11] (80.75%) by a 10% margin. On CIFAR-100, SAFLe (70.61%) again leads DeepAFL (66.98%) and AFL (58.56%), and on Tiny-ImageNet, our method (64.58%) continues to outperform DeepAFL (62.35%) and AFL (54.67%). This demonstrates the superior representational power of our proposed non-linear analytic architecture, which achieves higher accuracy than even the multi-round DeepAFL while maintaining single-round efficiency.
@@
Invariance Analysis Table 1 empirically shows the core theoretical benefit of our analytic formulation: invariance to data heterogeneity. While all iterative gradient-based methods suffer severe performance degradation as the data becomes more non-IID (e.g., $\alpha=0.01$ or $s=5$), our method does not. For instance, on CIFAR-100, FedAvg’s accuracy plummets from 56.62% ($\alpha=0.1$) to 32.99% ($\alpha=0.01$), a drop of over 23.6%. In stark contrast, all analytic methods are completely unaffected by the data distribution. Our proposed SAFLe achieves an identical 70.61% accuracy across all four non-IID settings on CIFAR-100, just as AFL (58.56%) and DeepAFL (66.98%) maintain their respective scores. This invariance also extends to the number of clients, as shown in Figure 5. While FedAvg’s performance clearly declines as $K$ increases from 100 to 1000, SAFLe’s accuracy remains constant, matching the invariance of AFL. This experimentally showcases SAFLe’s successful extension of the mathematical invariance of AFL to a more expressive non-linear model.
Communication Rounds Efficiency DeepAFL [25] scales AFL by adding depth (i.e., more layers), but at the direct trade-off of more communication rounds. As seen in Figures 4 and 2, DeepAFL must introduce two new communication rounds for every additional layer, requiring 41 rounds for a $T=20$ model. Our method, SAFLe, scales by adding width (i.e., more experts $E$), allowing us to scale the model’s expressivity while remaining a single-shot solution. SAFLe’s expressivity is decoupled from its number of communication rounds, which is always one round.
@@
Ablation: Bucketing Strategy The bucketing function used to discretize continuous backbone features is an important design choice. In Table 2, we compare three variants: Standard (Integer), One-Hot, and Binary Overlapping. Binary Overlapping performs best, while standard Integer quantization performs worst by a large margin. We believe this is because Integer quantization introduces a sharp “cliff effect,” where small changes across bin boundaries produce completely different indices and therefore very different embedding lookups, hurting generalization. One-Hot is slightly better but still assigns nearby inputs to disjoint representations. In contrast, Binary Overlapping preserves local similarity by giving nearby values similar bit patterns, allowing the model to learn more consistent embeddings and generalize better.
@@
Ablation: Different Backbones. To show that SAFLe’s gains are not tied to a specific backbone, we follow the AFL ablation protocol [11] and compare against AFL using the same three frozen pre-trained backbones: ResNet-18 [10], VGG11 [24], and ViT-B-16 [8]. DeepAFL [25] is not included here because it does not report this experiment and its code is not yet public. As shown in Table 3, SAFLe consistently outperforms AFL across all backbones and datasets. While stronger backbones improve both methods, SAFLe’s non-linear analytic head continues to provide a clear gain over AFL’s linear head, confirming that our advantage is robust to backbone choice.
Compared with recent one-shot federated methods built on strong pre-trained models, SAFLe remains superior. In particular, against TOFA [33], a recent training-free one-shot method based on a ViT-B/16 vision-language backbone, SAFLe achieves higher accuracy on both CIFAR-10 (95.31% vs. 93.18%) and CIFAR-100 (77.83% vs. 76.63%). This shows that SAFLe’s gains are not tied to a particular backbone family and remain robust even when compared against strong pre-trained one-shot baselines.
@@
Robustness to Label Noise. We also observe that SAFLe is reasonably robust to noisy labels, even though it does not use any explicit label-noise correction mechanism. On ISIC 2019 [26, 6, 12] with label noise rates $\rho=\{0.2,0.3,0.4,0.6\}$, SAFLe achieves 69.47%, 67.82%, 66.53%, and 62.94%, respectively. These results are competitive with the specialized noise-robust method FedNoRo [29], which reports 69.8%, 68.5%, 68.9%, and 63.3%, and consistently outperform FedCorr+LA [30], which reports 63.6%, 61.9%, 59.6%, and 53.9% at the same noise levels. This suggests that SAFLe inherits a useful degree of intrinsic robustness from its analytic formulation, even without task-specific noise handling.